# Consistency of Responses and Continuations Generated by Large Language Models on Social Media

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities in text generation, yet their emotional consistency and semantic coherence in social media contexts remain insufficiently understood. This study investigates how LLMs handle emotional content and maintain semantic relationships through continuation and response tasks using two open-source models: Gemma and Llama. By analyzing climate change discussions from Twitter and Reddit, we examine emotional transitions, intensity patterns, and semantic similarity between human-authored and LLM-generated content. Our findings reveal that while both models maintain high semantic coherence, they exhibit distinct emotional patterns: Gemma shows a tendency toward negative emotion amplification, particularly anger, while maintaining certain positive emotions like optimism. Llama demonstrates superior emotional preservation across a broader spectrum of affects. Both models systematically generate responses with attenuated emotional intensity compared to human-authored content and show a bias toward positive emotions in response tasks. Additionally, both models maintain strong semantic similarity with original texts, though performance varies between continuation and response tasks. These findings provide insights into LLMs’ emotional and semantic processing capabilities, with implications for their deployment in social media contexts and human-AI interaction design.
@@ Introduction
Previous investigations of emotional effects on social media have employed real-life digital experiments through content manipulation (Kramer, Guillory, and Hancock 2014). However, such methodologies raise ethical concerns regarding unauthorized manipulation of user content and have generated public discomfort (Boyd 2016). LLMs offer a more ethically sound approach to examining emotional dynamics in social contexts. Given their increasingly sophisticated human-like capabilities, LLMs are extensively employed in simulating social interactions (Gao et al. 2024a). This LLM-based simulation methodology presents two key advantages: (1) AI agents can serve as safer substitutes for human participants in extreme or sensitive scenarios, and (2) they enable more controlled experimental conditions, facilitating precise examination of relevant variables.
With the widespread adoption of LLMs in generating human-like content, it becomes imperative to understand the consistency of LLM-generated text and its potential societal impact. Accordingly, this study investigates LLM text generation tasks (response generation and continuation) through systematic analysis of emotional consistency and semantic similarity. By examining these dynamics within climate change communication—a highly polarized and emotionally charged domain—this research addresses the following questions:
@@
By answering the research questions, this study has the following contributions:
  • Although LLMs demonstrate remarkable human-like capabilities in text generation, the mechanisms underlying their outputs remain insufficiently understood. Through analysis of emotional consistency between human- and LLM-generated text in social media contexts (response and continuation tasks), this study provides a deeper, contextualized understanding of LLMs’ text generation performance. The identified distinctions between human- and LLM-generated content illuminate how these models navigate social media interactions, which prompts critical ethical considerations regarding AI’s role in human interaction.
  • Emotion represents a fundamental behavioral response and crucial element in social media discourse. While previous research has predominantly focused on human-human communication, raising ethical concerns (Ferrara and Yang 2015), this study implements an LLM-based simulation approach. This methodology replicates human-AI agent interactions while addressing ethical limitations inherent in traditional research approaches, offering an innovative and ethically sound framework for emotion research.
  • Through analysis of real-world social media datasets concerning controversial scientific topics such as climate change, this study simulates scenarios where LLM-enabled tools participate in public discourse, presenting both opportunities and challenges (Feng et al. 2024). These findings enhance our understanding of communication dynamics and catalyze discussions regarding LLMs’ role in emotional guidance within controversial scientific discourse on social media platforms.
@@
Figure 1 illustrates the overall setup of the experimental design. In our experiment, we used two open-source large language models: Gemma [^1] https://ai.google.dev/gemma and Llama [^2] https://ai.meta.com/llama/license/, developed by Google and Meta, respectively. Specifically, we chose Gemma2-27B-Instruct-Q8 and Llama3-70B-Instruct, respectively, both of which excel in both performance and robustness. We utilized Ollama [^3] https://ollama.com/ as a framework to enable the two open source models to run on our local server.
Ollama supports various functions such as model creation, content generation, chat and embedding calculation. In our study, we mainly used the chat and embedding calculation. For the response task, we called its chat function directly, and unlike normal content generation functions, it allows the large language model to talk directly to the input content, which means there is no limitation of prompt words. This enables the most intuitive observation of the state of the large language model as an interlocutor, yielding more realistic and direct data. When it comes to the continuation task, we used the content generation function, which needs a prompt to ask the LLMs to expand the text it received. Again, here we made our prompts as concise as possible to minimize the impact of the prompts on the model. We tell the LLMs that “Assuming you are the author of this text, stand in your shoes and continue to expand the passage as you understand it.”
Figure 1: Experimental pipeline of consistency evaluation for LLMs. Our experimental framework begins with human text input to two LLMs (Gemma and Llama), which perform two distinct tasks: continuation and response. The continuation task employs a specific prompt instructing the model to expand the text as its author, while the response task operates without explicit prompting to enable natural interaction. Following content generation, we implement emotion detection on the outputs, followed by comprehensive analyses. The framework concludes with parallel analyses of emotional content and semantic similarity to evaluate the consistency of LLM-generated content relative to the original human input.
@@
This study utilized climate change corpora collected from Twitter (now X) and Reddit. We collected data using the Twitter Search API by querying relevant keywords, including climate change”, climate science”, climate manipulation”, climate Engineering”, climate Hacking”, climate modification”, Global Warming”, carbon footprint”, and “The Paris Agreement”. For Reddit, we used data maintained by Pushshift from https://the-eye.eu/redarcs/. The Pushshift Reddit dataset consists of two sets of files: submissions and comments(Baumgartner et al. 2020). The same keywords were applied to filter Reddit data, and to compare the differences in emotions, we collected both posts and comments from both platforms.
@@
With the keywords, we obtained 5,768,822 Reddit comments and 76,596,654 tweets from Twitter. We used histograms to understand the basic distribution of data (Figure 2). To ensure temporal representation and minimize the impact of special events, we employed proportional sampling, selecting 200 rows per month systematically from Twitter data and 100 rows per month from Reddit data. This sampling strategy yielded a final dataset comprising 12,200 rows of Twitter data and 10,900 rows of Reddit data.
@@
In this study, we developed a methodology to analyze emotions in cross-platform social media data using a deep neural network-based model. We employed the twitter-roberta-base-emotion-multilabel-latest [^4] https://huggingface.co/cardiffnlp/twitter-roberta-base-emotion-multilabel-latest model from Hugging Face to examine the emotional content of both original texts and content generated by large language models. This model, built upon the RoBERTa-base architecture, is a fine-tuned version of “cardiffnlp/twitter-roberta-base-2022-154m” optimized on the SemEval 2018 Task 1 dataset—Affect in Tweets.
The twitter-roberta-base-emotion-multilabel-latest model identifies eleven distinct emotion categories: anticipation, joy, love, optimism, surprise, trust, anger, disgust, fear, pessimism, and sadness. The model outputs probability scores for each category, which serve as quantitative measures of emotional content for subsequent analysis. While previous studies used sentiment analysis (negative, positive, and neutral) for evaluating differences between human- and LLMs-generated text (Guo et al. 2023), our emotion-based approach provides a more granular and nuanced understanding of the underlying emotional states in posts.
Semantic Similarity
To assess both the comprehension capabilities of LLMs and their content generation accuracy, we employed cosine similarity as a quantitative measure of semantic alignment between LLMs’ outputs and human-generated texts. This methodology also enables the detection of semantic aberrations, commonly referred to as “hallucinations”, which frequently manifest in LLM outputs (Breum et al. 2023).
Cosine similarity, a fundamental metric in natural language processing, has demonstrated its utility across various text mining applications, including text classification, summarization, information retrieval, and question answering systems (Li and Han 2013). The mathematical foundation of this metric ensures robust comparison of textual similarities, making it particularly suitable for our analysis.
Due to the architectural differences between the two models, we implemented separate embedding calculations for LLM-generated and human-authored texts using both Gemma and Llama. The semantic similarity computations were performed using the cosine similarity function from the sklearn [^5] https://scikit-learn.org/stable/ library. For each platform, we derived four distinct similarity metrics quantifying the semantic relationships among four text pairs: Gemma’s continuation text, Gemma’s response text, Llama’s continuation text, and Llama’s response text, each compared against the original human text. To evaluate the comparative performance across different tasks and models, we employed the Mann-Whitney U-test to determine the statistical significance of variations in cosine similarity scores.
@@ Results
In this study, we examined the emotional transitions between human-generated text and LLM outputs in downstream tasks. We categorize 11 kind emotions into those that are positively oriented and those that are negatively oriented as followed(Robinson 2008) :
@@
Figure 3: Emotional Transition Analysis of LLM Response and Continuation Tasks in Reddit Comments. Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks, respectively. The y-axis represents source emotions from human text, while the x-axis indicates emotions in LLM-generated content. Cell values represent the proportion of emotional transitions between original and generated content. For example, in Figure 3a, the value 0.64 in the anger-to-anger cell indicates that 64% of originally angry texts maintained their emotional valence in Gemma’s continuation task. The intensity of each cell’s shading represents the proportion of emotional transition, with darker shades indicating higher transition frequencies.
Analysis of Figure 3a reveals that 62% of texts initially labeled as angry maintained their emotional valence in Gemma’s continuations. In contrast, only 11% of originally optimistic texts maintained their optimistic valence in the continuations. Other emotional categories, including anticipation, disgust, fear, joy, and sadness, predominantly shifted toward anger in the continuations, with transition rates of 29%, 44%, 30%, 39%, and 31%, respectively. Notably, optimism and surprise exhibited distinct patterns: 43% of optimistic texts preserved their emotional valence in the Gemma task, while texts expressing surprise demonstrated consistent emotional preservation. These findings suggest that the Gemma model exhibits a systematic tendency to transform diverse emotional expressions into anger during continuation tasks, indicating a bias toward negative emotional content, particularly anger. However, its performance with optimism and surprise demonstrates capacity for emotional preservation, suggesting selective ability to maintain certain emotional contexts throughout the generation process. In the analysis of emotional transitions during Gemma’s response task, we observed a significant shift in emotional valence, with more than 50% of initially angry texts transitioning toward anticipation and optimism in the responses.
Our analysis reveals that over 50% of texts with negative emotional valence transitioned toward positive expressions, particularly anticipation and optimism, in the response tasks. Texts with initial positive valence consistently preserved their affective characteristics, manifesting as anticipation, joy, or optimism. This pattern highlights Gemma’s systematic bias toward positive affect during response tasks. Significantly, a subset of original emotions still transitioned to anger in the responses, suggesting Gemma’s persistent sensitivity to anger-related content across both response and continuation tasks.
Examination of Llama 3’s performance, presented in Figure 3c and Figure 3d, demonstrates enhanced capability in emotional recognition and preservation during continuation tasks, maintaining original emotional valence with greater consistency.
The underlying mechanisms of emotional preservation in LLM-generated text appear to involve multiple factors: emotion-specific lexical choices, syntactic structures, and contextual integration during the generation process, coupled with the model’s capacity for sustaining emotional patterns. Emotions characterized by prominent semantic features (e.g., intense affective vocabulary) appear more readily preserved by the model.
Quantitative analysis of emotional transitions reveals that 21% of texts originally expressing anticipation, 48% expressing joy, and 17% expressing optimism transitioned to anger expressions. This pattern can be attributed to two primary factors: first, the semantic prominence of anger, characterized by explicit and intense features (e.g., critical, adversarial language) that are more readily recognized by the model during generation tasks; second, the contextual fragility of positive emotions, which typically require more complete contextual support, whereas LLM-generated content appears to favor more salient negative emotional expressions.
Consistent with Gemma’s response patterns, Llama’s response task demonstrated that, excluding fear, transitions to anticipation and optimism exceeded 50% across emotional categories. This consistent positive bias in response tasks across both models suggests a systematic emotional bias embedded during training, particularly evident in interactive contexts.
@@
Figure 4: Emotional Transition Analysis of LLM Response and Continuation Tasks in Twitter Comments. Panels a, b, c, and d illustrate emotional transitions in content generated by Gemma and Llama models during continuation and response tasks on Twitter, respectively. The y-axis represents the original emotions in human-authored tweets, while the x-axis shows the emotions detected in LLM-generated content. Each cell value represents the proportion of emotional transitions, with darker shades of red indicating higher transition frequencies. For example, cell values of $1.0$ would indicate complete preservation of the original emotion, while lower values indicate transitions to different emotional states. The heatmap reveals patterns in how each model preserves or transforms emotional content across response and continuation tasks, with notable variations in the handling of positive (e.g., joy, optimism) versus negative (e.g., anger, fear) emotions.
In addition to Reddit, we analyzed data from Twitter, which exhibits distinct discourse patterns surrounding climate change. Figure 4 illustrates the emotional transitions between original Twitter content and LLM-generated responses. In Gemma’s continuation and response tasks, we observed that regardless of the original emotional valence—whether positive (anticipation, joy) or negative (anger, disgust)—the generated content predominantly expressed anger and anticipation. Among these transitions, anger represented the highest proportion, while anxiety emerged as the predominant emotion, followed by anticipation. A notable distinction between the two tasks was the increased frequency of responses transitioning toward anticipation. In Llama’s continuation task, the original emotional content similarly demonstrated a predominant shift toward anger. In contrast, Llama’s response task exhibited a primary shift toward anticipation. The performance patterns of both Gemma and Llama models across continuation and response tasks reveal two key insights: first, Gemma’s heightened sensitivity to anger-related content, and second, the models’ systematic bias toward positive affect when functioning in interactive dialogue contexts.
@@
Figure 5: Emotional source analysis of LLM-generated content across platforms. Panels a, c, e, and g illustrate emotional transitions in Gemma and Llama models’ continuation and response tasks on Reddit data. Panels b, d, f, and h show corresponding results from Twitter data. Red lines indicate content generated from originally positive emotional text, while blue dashed lines represent content derived from negative emotional text. The absolute x-values represent the proportion of emotional source contributions in the generated content. The y-axis displays the emotional categories present in both original and generated content, while the x-axis ($-1.0$ to $1.0$) indicates the strength and direction of emotional transitions.
We analyze the emotional sources of LLM-generated content by examining the relationship between input and output emotions. In Gemma’s continuation task, Figure 5a reveals that positive emotions in generated content primarily derive from positive emotional sources. For instance, joy-labeled content originates from anticipation (32.04%), joy (25.24%), and optimism (13.59%). Conversely, content expressing negative emotions predominantly stems from negative emotional sources, with anger-labeled content derived primarily from original anger (61.8%) and disgust (13.03%). This pattern suggests Gemma’s tendency to maintain emotional valence consistency during content continuation.
Figure 5c illustrates Llama’s continuation task results, which differ from Gemma’s pattern. Most emotional content, except for optimism and joy, originates predominantly from negative emotional sources. While the preservation of negative emotional sources aligns with Gemma’s behavior, Llama distinctively generates positive emotional content primarily from negative emotional sources, potentially indicating an inherent positive affect bias.
Figure 5b,d illustrate the response tasks for Gemma and Llama, respectively. In Gemma’s responses, most positive emotional content, except joy, originates from predominantly negative emotional sources. Negative emotional content maintains its source valence, with anger-labeled content derived 69.55% from original anger expressions. Similarly, Llama’s responses demonstrate consistent transformation of negative emotions into positive ones, exemplified by anticipation-labeled content originating 37.23% from anger. However, negative emotional content maintains its original valence.
These findings demonstrate LLMs’ systematic transformation of negative emotions into positive ones during response tasks, while simultaneously exhibiting some degree of negative emotional preservation in their responses.
@@
Beyond examining emotional transitions, we investigated the quantitative differences in emotional intensity between LLM-generated and human-authored content. This analysis specifically focused on determining whether LLM-generated content exhibits higher or lower emotional intensity compared to human expressions. To quantify emotional content, we employed a probabilistic model that assigns normalized scores (0 to 1) to each emotional category within the text. These probability values were interpreted as emotional intensity scores (Miyazaki et al. 2024). We categorized emotional expressions into five distinct intensity groups based on these scores. Statistical analysis comprised an ANOVA test to evaluate differences in emotional intensity across groups, followed by Tukey’s post-hoc test to identify significant pairwise variations(Elnaggar, Mohamed, and Gehan 2024).
Table 1: Test for significant differences in affective values between groups

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable capabilities in text generation, yet their emotional consistency and semantic coherence in social media contexts remain insufficiently understood. This study investigates how LLMs handle emotional content and maintain semantic relationships through continuation and response tasks using three open-source models: Gemma, Llama3 and Llama3.3 and one commercial Model:Claude. By analyzing climate change discussions from Twitter and Reddit, we examine emotional transitions, intensity patterns, and semantic consistency between human-authored and LLM-generated content. Our findings reveal that while both models maintain high semantic coherence, they exhibit distinct emotional patterns: these models show a strong tendency to moderate negative emotions. When the input text carries negative emotions such as anger, disgust, fear, or sadness, LLM tends to generate content with more neutral emotions, or even convert them into positive emotions such as joy or surprise. At the same time, we compared the LLM-generated content with human-authored content. The four models systematically generated responses with reduced emotional intensity and showed a preference for neutral rational emotions in the response task. In addition, these models all maintained a high semantic similarity with the original text, although their performance in the continuation task and the response task was different. These findings provide deep insights into the emotion and semantic processing capabilities of LLM, which are of great significance for its deployment in social media environments and human-computer interaction design.
@@ Introduction
Previous investigations of emotional effects on social media have employed real-life experiments through content manipulation (Kramer, Guillory, and Hancock 2014). However, it raise ethical concerns regarding manipulation of user content and public discomfort (Boyd 2016). Recent work demonstrates that LLM-based simulations can be a useful tool for modeling social interactions (Gao et al. 2024a): (1) AI agents can serve as safer substitutes for human participants in extreme or sensitive scenarios, and (2) they enable more controlled experimental conditions, facilitating precise examination of relevant variables.
With the widespread adoption of LLMs in generating human-like content, it becomes imperative to understand the consistency of LLM-generated text and its potential societal impact. Moreover, with a growing variety of LLMs being adopted in real-world settings, there remains a lack of systematic, comparative evaluations of how consistent/different these models perform.
Accordingly, this study investigates LLM text generation tasks (response generation and continuation) through systematic analysis of emotional consistency and semantic similarity. By examining these dynamics within climate change communication—a highly polarized and emotionally charged domain—this research addresses the following questions:
@@
By answering the research questions, this study has the following findings and contributions:
  • We systematically evaluated emotional and semantic consistency between LLM-generated and human-authored social media texts, finding that (1) LLMs’ emotional expressions differ significantly from original texts across both continuation and response tasks, and across different model types; (2) LLMs maintain high semantic similarity with original content, indicating strong capabilities in understanding and generating human-like text.
  • We found that all models exhibit significantly lower emotional intensity in both tasks, suggesting that LLMs may struggle to convey the full emotional depth of human-authored content.
  • We discussed the real-world implications of these emotional differences and semantic proximity — including how LLMs may influence emotional engagement, serve as tools for moderating polarized debates, or, conversely, be misused for emotional manipulation.
  • We proposed an ethically grounded simulation framework using LLM agents to explore emotional dynamics around controversial topics on social media, avoiding privacy and consent issues inherent in traditional research (Ferrara and Yang 2015). We also adopted the “LLM-as-judge” approach to assess semantic similarity, reducing reliance on time-consuming and costly human annotations.
@@
Figure 1 illustrates the overall setup of the experimental design. Our core experiment involves using LLMs to interact with human-generated text data, which consists of social media posts collected from real users. In our experiment, we employed three open-source large language models—Gemma [^1] https://ai.google.dev/gemma, Llama 3222https://ai.meta.com/llama/license/, and Llama 3.3—developed by Google and Meta, respectively, as well as one commercial model, Claude 3.5, developed by Anthropic [^3] https://www.anthropic.com/claude. Specifically, we utilized the following model variants: Gemma2-27B-Instruct-Q8, Llama3-70B-Instruct, Llama3.3-70B, and Claude-3-5-Haiku-20241022-X. These models were selected due to their strong performance and robustness. We utilized Ollama [^4] https://ollama.com/ as a framework to enable the two open source models to run on our local server.
In our study, we mainly used the chat of Ollama. For the response task, we directly used the chat interface, which allows LLMs to interact with the input content without requiring explicitly crafted instruction prompts. This setup offers a more natural and direct observation of the LLM’s behavior as a conversational agent. However, we acknowledge the reviewer’s point and clarify that although we did not provide any additional instruction prompt, the original human post itself serves as the input prompt to the model. For the continuation task, we used the content generation function, where concise prompts were employed to minimize their influence on the model, allowing for a clearer focus on the model’s response.We tell the LLMs that “Assuming you are the author of this text, stand in your shoes and continue to expand the passage as you understand it.”
Figure 1: Experimental pipeline of consistency evaluation for LLMs. Our experimental framework begins with human text input to four LLMs , which perform two distinct tasks: continuation and response. The continuation task employs a specific prompt instructing the model to expand the text as its author, while the response task operates without explicit prompting to enable natural interaction. Following content generation, we implement emotion detection on the outputs, followed by comprehensive analyses. The framework concludes with parallel analyses of emotional content and semantic consistency to evaluate the consistency of LLM-generated content relative to the original human input.
@@
This study utilized climate change corpora collected from Twitter (now X) and Reddit. We collected data using the Twitter Search API by querying relevant keywords, including “climate change”, “climate science”, “climate manipulation”, “climate Engineering”, “climate Hacking”, “climate modification”, “Global Warming”, “carbon footprint”, and “The Paris Agreement”. For Reddit, we used data maintained by Pushshift from https://the-eye.eu/redarcs/. The Pushshift Reddit dataset consists of two sets of files: submissions and comments  (Baumgartner et al. 2020). The same keywords were applied to filter Reddit data, and to compare the differences in emotions, we collected both posts and comments from both platforms.
@@
With the keywords, we obtained 5,768,822 Reddit comments and 76,596,654 tweets from Twitter. We used histograms to understand the basic distribution of data (Figure 2). Twitter and Reddit are two major social media platforms that differ from each other. Twitter is designed to be open, concise, and immediate through short posts and real-time updates, while Reddit is a community-based medium designed for deep, deliberative, and topical discussions(Parsa et al. 2022; Treen et al. 2022a). These differences in platform design not only shape how content is generated and shared, but also attract different user groups with different interests and engagement styles, which provides a valuable resource for understanding different modes of public discussion(Ruan et al. 2022).
We support these distinctions with some quantitative evidence using the average word length, word cloud. The results show that the average length of Reddit text is 656.66 words, while the average length of Twitter text is 246.34 words. This discrepancy reflects fundamental differences in how the two platforms are used. Reddit is structured around topic-based discussions within subreddits, encouraging longer, more detailed posts that often resemble mini-essays or narratives. The word cloud chart (see the appendixLABEL:wordcloud) shows that the expressions of users on the two platforms under the same topic are also different. For example, more people on Reddit are discussing ”climate change”, while more people on Twitter use ”global warming”.
To construct datasets representative of the discussion dynamics surrounding climate change over an extended period and to mitigate potential biases introduced by sudden climate change events, we employed a time-stratified sampling approach on the raw data. Specifically, based on the collected corpus of 5,768,822 Reddit comments and 76,596,654 Twitter tweets, we initiated the process by partitioning the data by month. Subsequently, within each monthly segment, we conducted systematic sampling, extracting 200 records from the Twitter data and 100 records from the Reddit data. This method of monthly proportional sampling was implemented to ensure a uniform distribution of the dataset across the temporal dimension, thereby reducing the influence of transient extreme events or fluctuations in topic popularity during specific timeframes. This strategy aims to enable the final dataset to more comprehensively reflect the content and evolving trends of discussions pertaining to climate change across different temporal stages. Ultimately, this process yielded an analytical dataset comprising 12,200 Twitter data points and 10,900 Reddit data points.
@@
In this study, we developed a methodology to analyze emotions in cross-platform social media data using a deep neural network-based model. We employed thej-hartmann/emotion-english-distilroberta-base [^5] https://huggingface.co/j-hartmann/emotion-english-distilroberta-base model from Hugging Face to examine the emotional content of both original texts and content generated by large language models. This model, built upon the RoBERTa-base architecture, is a fine-tuned checkpoint of “DistilRoBERTa-base” the datasets contain emotion labels for texts from Twitter, Reddit, student self-reports, and utterances from TV dialogues.
The emotion-english-distilroberta-base model identifies seven distinct emotion categories:joy, surprise, neutral, anger, disgust, fear and sadness. The model outputs probability scores for each category, which serve as quantitative measures of emotional content for subsequent analysis. While previous studies used sentiment analysis (negative, positive, and neutral) for evaluating differences between human and LLMs-generated text (Guo et al. 2023), our emotion-based approach provides a more granular and nuanced understanding of the underlying emotional states in posts.
Semantic Consistency
We design a evaluation framework that uses a rubric-guided expert-like LLM to score consistency and faithfulness. Here we hired the model:gemini-2.0-flash-thinking-exp-1219 as our expert LLM. This allows us to directly assess whether a model-generated response is logically aligned with the user’s comment, and please see the appendix for specific prompt designLABEL:lst:build-prompt. After obtaining the evaluation results of the LLM-as-judge, the results are reviewed by a human evaluator. Human evaluator judges whether the evaluation results of the LLM are reasonable and accurate. In the end, the consistency rate between the two is maintained at 75%.
Automatic evaluations of instruction following abilities in LLMs has recently received significant attention (Zheng et al. 2023). Given the significant time and cognitive effort required for human evaluation of large-scale generated content, we adopted a stratified random sampling approach to ensure representativeness across different model-platform-task combinations. Our full dataset consists of different sub-corpora, each corresponding to a specific combination of model (e.g., Gemma, Llama 3, Llama 3.3, Claude), platform (Twitter, Reddit), and task type (response or continuation), with approximately 10,000 instances per sub-corpus. To balance coverage and feasibility, we randomly sampled 30 instances from each sub-corpus, resulting in a total of 480 samples for evaluation.
Manual expert evaluation is costly. This hybrid evaluation strategy combines manual judgment with the assistance of a LLM. Although the sample size is small, it covers diversity across platforms, tasks, and models. We acknowledge in the limitations section that this small sample size may affect the generalizability of our findings.
@@ Results
In this study, we examined the emotional transitions between human-generated text and LLM outputs in downstream tasks. We categorize 7 emotions as positive emotionally oriented and negatively oriented as well as neutral emotions as defined by itself, as follows (Robinson 2008) :
@@
Figure 3: Emotional Transition Analysis of LLM Response and Continuation Tasks in Reddit Comments. Panels a, b, c, d, e, f, g, and h illustrate emotional transitions in content generated by Gemma, Llama and Claude models during continuation and response tasks, respectively. The y-axis represents source emotions from human text, while the x-axis indicates emotions in LLM-generated content. Cell values represent the proportion of emotional transitions between original and generated content. For example, in Figure 3a, the value 0.34 in the anger-to-anger cell indicates that 34% of originally angry texts maintained their emotional valence in Gemma’s continuation task. The intensity of each cell’s shading represents the proportion of emotional transition, with darker shades indicating higher transition frequencies.
Analysis of Figure 3a reveals that in Gemma’s continuation tasks, 34% of the texts initially labeled as ’anger’ maintained their emotional, and 38% of the texts originally categorized as ’fear’ also preserved the ’fear’ emotion in the continuation task. For other emotion categories, including ’disgust,’ ’sadness,’ ’neutral,’ ’joy,’ and ’surprise,’ the proportions of texts that retained the same emotional label were 22%, 13%, 40%, 10%, and 1%, respectively, demonstrating a capability for emotion preservation.
It is noteworthy that for all original emotion labels, a considerable proportion of the texts shifted to a neutral label in the continuation task. Concurrently, a significant portion of texts also transitioned to the ’anger’ label during the continuation. These findings suggest that the Gemma model possesses a certain ability to recognize and perpetuate the original emotions. In subsequent tasks, it exhibits a systematic tendency to convert various emotional expressions into a neutral sentiment. The results in Figure 3a also, to some extent, indicate the model’s sensitivity to negative emotions in the continuation task, particularly anger.
In analyzing the emotional shifts within Gemma’s response tasks, we observed significant changes in emotional conversion. For texts with originally positive emotions such as ’joy’ and ’surprise,’ as well as those with negative emotions like ’anger,’ ’disgust,’ ’fear,’ and ’sadness,’ over 50% of their emotion labels were converted to neutral in the response task. Our analysis reveals a systematic bias in Gemma towards rational, neutral sentiment in these response tasks. Notably, a portion of the original emotions still transformed into ’anger’ in the responses; for instance, ’disgust’ showed a 21% conversion rate to ’anger,’ while ’surprise’ had a 14% conversion rate. This suggests Gemma’s sustained sensitivity to anger-related content across both response and continuation tasks.
Figure 3c and Figure 3d illustrate the performance of Llama 3, showing that it has the same ability to recognize and preserve emotions in continuation and response tasks, and is able to maintain the original emotional valence more consistently. Similarly, we see a neutral emotion bias in the Llama3 model.
In the response and continuation tasks of the commercial model Claude, we observed that more original text emotions were more neutral in downstream tasks. At the same time, in the same model family, we observed that the Llama 3.3 model was also more inclined to generate neutral content in response and continuation than Llama3.
Beyond Reddit, our analysis also included Twitter data, which revealed distinct discourse patterns surrounding climate change topic. Figure 4 illustrates the emotion shift between the original Twitter content and the replies generated by the LLM. The emotion performance patterns of the Gemma, Llama and Claude models on Twitter texts in continuation and response tasks yielded two key insights: Firstly, Gemma demonstrated heightened sensitivity to content related to anger; secondly, for models exhibited a systematic bias towards neutral sentiment when operating in interactive conversational contexts. third, both Claude and Llama3.3 models tend to generate content with neutral emotions. In addition, it is worth noting that in the continuation task of Llama3.3, its ability to continue emotions is better than Llama3.
@@
Figure 4: Emotional Transition Analysis of LLM Response and Continuation Tasks in Twitter Comments. Panels a, b, c, d e, f, g, and h illustrate emotional transitions in content generated by Gemma, Llama and Claude models during continuation and response tasks on Twitter, respectively. The y-axis represents the original emotions in human-authored tweets, while the x-axis shows the emotions detected in LLM-generated content. Each cell value represents the proportion of emotional transitions, with darker shades of red indicating higher transition frequencies.
@@
Figure 5: Emotional source analysis of LLM-generated content across platforms. Panels an illustrate emotional transitions in Gemma, Llama and Claude models’ continuation and response tasks on Reddit and Twitter data. The red bars indicate that the original text corresponding to the generated text is positive emotions such as joy and surprise, the blue bars indicate that the original text is negative emotions such as anger, disgust, fear and sadness, and the green bars indicate that it comes from text with neutral emotions. The y-axis displays the emotional categories present in both original and generated content.
We analyzed the emotional sources of LLM-generated content by examining the relationship between input and output emotions. As shown in Figure 5a, the results show that Gemma did not simply copy or continue the original emotion of the input text, but perform significant emotion modulation. A core finding is that the model has a general tendency to moderate negative initial emotion. Specifically, when the original text carries negative emotion (as shown by the blue bar in the figure 5, representing anger, disgust, fear, or sadness), a considerable proportion of the generated content emotion will turn into neutral (source shown by the green bar) or even positive emotion (source shown by the red bar, representing joy or surprise). For example, in the continuation and response tasks of multiple models (such as Gemma, Llama, Claude), as well as on the Reddit and Twitter datasets, it can be observed that the generated text with ”neutral” or ”joy” emotion has a significant negative emotion contribution in its original emotion input. In addition, the model also shows a general trend towards neutral emotion, that is, the generated neutral emotion text comes from a wide range of original positive, negative, and neutral inputs. Even if the original input is positive, the generated content has a certain probability of being adjusted to neutral. These emotion conversion patterns show high consistency between different LLMs, task types (text continuation and response), and data sources (Reddit and Twitter), suggesting that this may be an inherent characteristic or common training result of current mainstream LLMs in emotion processing.
@@
We analyzed the differences in emotional intensity between LLM-generated and human-authored content, focusing on whether LLMs exhibit higher or lower emotional intensity. Emotional content was quantified using a probabilistic model assigning normalized scores (0 to 1) to each emotional category, interpreted as intensity scores (Miyazaki et al. 2024). These scores were grouped into five intensity levels. Statistical analysis included ANOVA to compare group differences and Tukey’s post-hoc test for significant pairwise variations(Elnaggar, Mohamed, and Gehan 2024).
Table 1: ANOVA test of different model emotions